OpenStudy (anonymous):

Which of the following ratios is not equivalent to 6:10? a. 3/5 b. 9/15 c. 48/80 d. 24/45

OpenStudy (anonymous):

Just divide the numerator by the denominator. A. 3 / 5 = 0.6 B. 9 / 15 = 0.6 C. 48 / 80 = 0.6 d. 24 / 45 = 0.533, there's your answer.

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

Need anything else?

OpenStudy (anonymous):

Yupp. Keep your eyes open in the open questions section.